OpenStudy (anonymous):

what is 2.4 and 7.4 squared

OpenStudy (whpalmer4):

do you mean \[2.4^2+7.4^2\]or\[(2.4+7.4)^2\]

OpenStudy (anonymous):

how do u do the squared thing

OpenStudy (anonymous):

\[a^{2}+2.4^{2}=7.4^{2}\]

OpenStudy (anonymous):

try times it by it self once but here is 2.4 (5.76) and 7.4 (54.76)

OpenStudy (whpalmer4):

\[a^2+2.4^2 = 7.4^2\]\[a^2=7.4^2-2.4^2\]\[a^2 = 54.76-5.76 =\]then take the square root to find the value of \(a\)

OpenStudy (anonymous):

thx